A company will begin stocking remote control devices. Expected monthly demand is 700 units. The controllers will be purchased from RCD supplier offering price breaks for quantities. RCD’s price list is as follows: Quantity Range Unit Price 0 to 299 16.00 300 to 399 15.80 400 to ? 15.10 Given that ordering cost is $40 and annual holding cost is 0.25 of unit price per unit, provide the following results: • What is the feasible quantity? o What is the total annual ordering cost at this quantity? $ o What is the total annual holding cost at this quantity? $ o What is the total annual material cost at this quantity? $ o What is the total annual cost at this quantity? $ • What is the most economical quantity to order? o What is the savings in total annual cost at this quantity? $ FORMAT FOR ENTRIES: • All quantities and costs are in two decimals after the dot. (8457.20, for example) • DO NOT ENTER $ sign for costs • If no saving in costs is available ENTER 0.00
Answer #1

Expected demand 700 units per month or 700*12=8400 annually

Pricing=P

0-299 : 16=P0

300-399 : 15.8=P1

400 above : 15.1=P2

Ordering Cost= Co=40

Holding cost=Ch=0.25 per unit price=0.25*P

For P=P0=16

Feasible quantity=Economic order quantity (EOQ)= squareroot(2*Demand*Ordering Cost/Holding Cost)

=sqrt(2*8400*40/(0.25*16))= sqrt(168000)= 410 units approx

For p=P1=15.8

EOQ=sqrt(2*8400*40/(0.25*15.8))=412.46 units

For P=P2=15.1

EOQ=sqrt(2*8400*40/(0.25*15.1))=421.91 Units

All other costs remaining same 421.91 units will give lowest cost, as it will have the lowest reorder cost due to high volume.